OpenStudy (anonymous):

Factor completely: 3x2 + 2x - 1 Answer (3x + 1)(x + 1) (3x + 1)(x - 1) (3x - 1)(x + 1) (3x - 1)(x - 1)

OpenStudy (anonymous):

(x+1) (3 x-1)

OpenStudy (anonymous):

you should explain the answers

OpenStudy (anonymous):

if you feel the need to explain, then by all means explain

OpenStudy (anonymous):

you simply foil it

OpenStudy (anonymous):

try expanding each of the options eg (3x + 1)(x - 1) = 3x(x - 1) + 1(x - 1) = 3x^2 - 3x + x - 1 = x^2 - 2x - 1 - its not this one